OpenStudy (anonymous):

what are the roots of x^2-3x+7=0

OpenStudy (anonymous):

Use the quadratic equation -b +- sqrt(b^2 .....

OpenStudy (ybarrap):

x^2-3 x+7 = 0 Subtract 7 from both sides: x^2-3 x = -7 Add 9/4 to both sides: x^2-3 x+9/4 = -19/4 Factor the left hand side: (x-3/2)^2 = -19/4 Take the square root of both sides: abs(x-3/2) = (i sqrt(19))/2 Eliminate the absolute value: x-3/2 = -(i sqrt(19))/2 or x-3/2 = (i sqrt(19))/2 Add 3/2 to both sides: x = 1/2 (3-i sqrt(19)) or x-3/2 = (i sqrt(19))/2 Add 3/2 to both sides: x = 1/2 (3-i sqrt(19)) or x = 1/2 (3+i sqrt(19))